OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza? twelve point five six square inches forty-seven point one zero square inches eighty-seven point nine two square inches three hundred fifty-one point six eight square inches

OpenStudy (anonymous):

Shouldn't it be pizzas i mean who orders pie at a pizza shop anyway, first find the area of both the large and small pizza with the formula: π(r^2) radius is half the diameter. then find the difference ignore the medium.

OpenStudy (aum):

Difference in area = \(\Large \pi (r_2^2 - r_1^2) = \pi (r_2+r_1)(r_2-r_1)\)

OpenStudy (anonymous):

Mr. Trundle used exactly one hundred twenty-eight ceramic tiles to cover the floor of a shower. If each tile is a square with 6-inch sides, what is the area of the shower floor? seven hundred sixty-eight square inches one thousand five hundred thirty-six square inches three thousand seventy-two square inches four thousand six hundred eight square inches

OpenStudy (anonymous):

okay, what you need is the area of 128 6-inch tiles, right? so 6x6=36 < this is area of one tile 128x36=? <the total area